Sam is planning to open a dockless scooter company in Fayetteville and wants to estimate the amount of time a typical user would rent one. She visits a similar college town to investigate, where she randomly samples 20 customers and asks each to report the number of minutes the scooter was used. The average rental time was 142.21 minutes. Sam knows from national data that scooter rental time is a normally distributed variable with a standard deviation of 26.09 minutes.

Sam would like to use the 90% confidence level; however, Sam’s business partner prefers to use a lower confidence level. Which of the following z or t values would be used to calculate an 80% confidence interval?

Group of answer choices

a. t = 1.33

b. z = .84

c. t = .69

d. z = 1.28
Answer #1

Given CI level is 80%, hence α = 1 - 0.8 = 0.2
α/2 = 0.2/2 = 0.1, Zc = Z(α/2) = 1.28

option D

z = 1.28